pm (3.00, -4.00) m and (700, 7.00) m 14.9
Answer #1

Given : r = (6.30 pm 0.30) cm , m = (1.71 pm 0.06) kg

Solution:

Step:1 Determine density

We know that density is given by :

Density = mass / volume

Volume of a sphere , V = (4/3)pir3

i.e. ho = m/V

= 1.71 kg / [4/3×3.14×(6.30×10-2)3]

= 1.63 × 103 kg/m3

Step:2 Determine uncertainty

Fractional error in the density is given by:

7m

= 0.06/1.71 + 3(0.30)/6.30

= 0.178

Total absolute uncertainty is given by:

(0.178)*(1.63×103) = 0.29×103

Step:3 Round and standard form

hopmΔρ =  (1.63 pm 0.29)×103 kg/m3